# Explaining word embeddings with perfect fidelity: A case study in predicting research impact

### v2

@@ Introduction
This section describes the workflow of our methodology for predicting the impactfulness of words as well as entire documents. A reliable prediction of research impact for any research paper only based on its abstract is a difficult task. In our research, we address a narrower subproblem of predicting which research will be groundbreaking, that is, we are interested in identifying the most influential research – both papers and newly emerged entities, such as drugs or virus strains in the CORD-19 training data context. The focus on the most confident predictions makes the research impact and explainability problem more tractable.
@@
The most significant property of embeddings for this study is that when words $w,w^{\prime}\in\mathcal{W}$ frequently appear “close” to each other in the training dataset $\mathcal{A}$, then $D(\widetilde{w},\widetilde{w}^{\prime})$ is “small”, where $D$ stands for some distance measure on $\mathbb{R}^{d}$, such as the cosine similarity or Euclidean distance.
@@
$\textit{SCORE}(A):=L(\widehat{\beta}_{0}+\widehat{\beta}^{T}\widetilde{A}).$ (4)
@@
$\textit{SCORE}(w)=L(\widehat{\beta}_{0}+\widehat{\beta}^{T}\widetilde{w}).$ (5)
@@
The data utilized for this study is derived from the CORD-19 corpus [11], which consists primarily of research papers pertaining to biology and medicine relevant to SARS-CoV-2 and that were published between 1952-2023. Specifically, we employed the version of the corpus downloaded on April 23, 2022, comprising a total of 992,921 papers. Paper citation counts were extracted from Opencitations.org. For the training of classifiers, we used only paper abstracts from the years 2016-2021 for which the Opencitation data were available.
@@
$\textit{CIT}(i,t)$ $:=\text{number of citations of paper $A_{i}$ originated in year $t$},$ (7)
@@
$y_{i}$ =$=\left\{\begin{array}[]{ll}1,&\text{if $\textit{CIT}(i,\tau)>\text{median}_{1\leq k\leq n}\{\textit{CIT}(k,\tau)\mid\textit{PUB}(k)\in T\}$,}\\ 0&\text{otherwise,}\\ \end{array}\right.$ (11)
where $\hskip 142.26378pt\mbox{where }i\in\{\iota\mid 1\leq\iota\leq n,\textit{PUB}(\iota)\in T\}.$ (12)
@@
Publication Year low high low high low high low high low high low high
@@
For training, we use 80% of available papers from publication years $t-1,t-2,\dots$ with the target being derived from citation counts of these papers accumulated by year $t-1$. The test papers were those published in year $t$ with the target being derived from citation counts also from year $t$. This is a particularly challenging setup because the prediction of citations for papers published in the same year is often very difficult.
@@
Since the last year needs to be used for testing, models trained under time-split testing are not trained on the latest papers, which negatively affects their ability to evaluate the impact of recent research. This semi-split evaluation technique best corresponds to how we envisage that the models would be used by leveraging the latest available citation counts for testing. Let $t$ be the most recent publication year. For training, we take 80% papers from $t-1$ to $t$, using citation counts accumulated up to the end of year $t$. The remaining 20% of the $t$-cohort form the test set and are evaluated by the citations they collect during the following calendar year $t+1$. Thus, the model learns from the freshest vocabulary while still being judged on genuine one-year-ahead citation predictions.
@@ Results
Although in terms of random-split evaluation, the BoW method most often has a higher AUC , LR-W2v model outperforms LR-BoW in most of the time-split & semi-split scenarios ($\{3\}$,$\{6\}$,$\{12\}$). The probable reason why in scenario $\bf\{9\}$, both models are failing as AUC is close to 0.5, is that the training occurred on pre-pandemic data and papers published during the pandemic were used for testing. The interest of researchers (and consequently the citations) shifted considerably as a result of the pandemic.
@@
where ##\## denotes the number of elements of a set and $H_{2021}$ is the set of highly-cited papers in 2021 computed as follows:
$H_{2021}$ =$=\{i\in\{1,\dots,n\}\mid PUB(i)\in\{2019,2020\},$
@@
$\textit{NumArt}(w)=\#\{i\in\{1,\dots,n\}\mid w\in A_{i},\ 2020\leq\textit{PUB}(i)\leq 2022\}\}.$ (16)
@@
The quality of the journal is measured through the ISI impact factor, which is computed as the citation counts to papers the journal published in 2019-2020 from papers published in the year 2021 in other journals with an impact factor divided by the number of papers published by the journal in the years 2019-2020. As a reference for a high-impact prediction, we use the median impact factor among the top 10% journals in the category “Health sciences” category (OECD category FORD 3030, data for the year 2021). To represent the research reputation of the authors, we use the maximum of the first and last authors’ Hirsch-index (H-index) values. The H-index is defined as the number of papers published by the author that have at least as many citations as this number. As a reference for a high-impact prediction, we use the median H-index of full professors, which was determined by analyzing the medical education journal editorial boards [26]. Previous research has found that the H-index was a good indicator of citation in the first few years after publication [25].
We also include additional impact metrics from social media as retrieved from Altmetrics.com in September 2022: the number of tweets mentioning the research and the aggregate attention received on social media (such as the number of Facebook likes) weighted by the relative reach of each type of source (Facebook, YouTube, Twitter).
@@
Figure 4: This figure shows the top papers by predicted impact (Score) published in 2021. The reference impact factor (blue line) represents the median among the top 10% journals in the category “Health sciences” category. The reference H-index (red line) corresponds to the median H-index of full professors. The attention score and tweets were retrieved from Altmetrics.com.
@@ Discussion
In Table 1(d), we also considered similar and important features related to the drug remdevisir, which has been used as a diagnostic word to assess word similarities for COVID-19 literature. In this case, all of the top important features were small-molecule COVID-19 drugs, with one exception being the antibody drug sarilumab approved for rheumatoid arthritis and repurposed for COVID-19. We note again that the suffix -ir, indicating an antiviral compound, is found for 8 of the 14 most important words similar to remdesivir, suggesting that subtoken approaches may provide additional benefits. Included in these top hits are some synonyms, e.g. molnupiravir and eidd-2801 (a preclinical name for this drug), lopinavir-ritonavir as well as lopinavir/ritonavir and lpv/r. There are 11/14 instances of first mentions in the literature prior to 2020, all representing drugs that were repurposed for COVID-19. Thus, we note that with our methodology, it is also possible to automatically identify instances of drug repurposing from the literature.
@@ Conclusions
The work of M. Černý was supported by Czech Science Foundation under project 23-06139S. The work of L. Beranová Dvořáčková was supported by grant IGA 16/2022 “PRECOG: Predicting REsearch COncepts of siGnificance”. T. Kliegr was supported by the long-term support for research activities of FIS VSE and the EU under agreement No 857446 (HeartBit 4.0). The authors would like to thank Karel Douda for acquiring the citation counts.

### v3

@@ Introduction
This section describes the workflow of our methodology for predicting the impactfulness of words as well as entire documents. A reliable prediction of research impact for any research paper only based on its abstract is a difficult task. In our research, we address a narrower subproblem of predicting which research will be groundbreaking. That is, we are interested in identifying the most influential research – both papers and newly emerged entities, such as drugs or virus strains in the CORD-19 training data context. The focus on the most confident predictions makes the research impact and explainability problem more tractable.
@@
The most significant property of embeddings for this study is that when words $w,w^{\prime}\in\mathcal{W}$ frequently appear “close” to each other in the training dataset $\mathcal{A}$, then D(w~,w~)D($\widetilde$$\widetilde{w}^{\prime}$) is “small”, where $D$ stands for some distance measure on $\mathbb{R}^{d}$, such as the cosine similarity or Euclidean distance.
@@
[…]
@@
[…]
@@
The data utilized for this study is derived from the CORD-19 corpus [11], which consists primarily of research papers pertaining to biology and medicine relevant to SARS-CoV-2 and that were published between 1952-2022. Specifically, we employed the version of the corpus downloaded on April 23, 2022, comprising a total of 992,921 papers. Paper citation counts were extracted from Opencitations.org. For the training of classifiers, we used only paper abstracts from the years 2016-2020 for which the Opencitation data were available.
@@
[…]
@@
[…]
$\hskip 142.26378pt\mbox{where }i\in\{\iota\mid 1\leq\iota\leq n,\textit{PUB}(\iota)\in T\}.$ (12)
@@
Publication
Year
low high low high low high low high low high low high
@@
For training, we use 80% of available papers from publication years $t-1,t-2,\dots$ with the target being derived from citation counts of these papers accumulated by year t1t-1. The test papers were those published in year $t$ with the target being derived from citation counts also from year $t$. This is a particularly challenging setup because the prediction of citations for papers published in the same year is often very difficult.
@@
Since the last year needs to be used for testing, models trained under time-split testing are not trained on the latest papers, which negatively affects their ability to evaluate the impact of recent research. This semi-split evaluation technique best corresponds to how we envisage that the models would be used by leveraging the latest available citation counts for testing. Let $t$ be the most recent publication year. For training, we take 80% papers from t1t-1 to $t$, using citation counts accumulated up to the end of year $t$. The remaining 20% of the $t$-cohort form the test set and are evaluated by the citations they collect during the following calendar year $t+1$. Thus, the model learns from the freshest vocabulary while still being judged on genuine one-year-ahead citation predictions.
@@ Results
Although in terms of random-split evaluation, the BoW method most often has a higher AUC , LR-W2v model outperforms LR-BoW in most of the time-split & semi-split scenarios ($\{3\}$$\{6\}$$\{12\}$). The probable reason why in scenario $\bf\{9\}$, both models are failing as AUC is close to 0.5, is that the training occurred on pre-pandemic data and papers published during the pandemic were used for testing. The interest of researchers (and consequently the citations) shifted considerably as a result of the pandemic.
@@
where #\# denotes the number of elements of a set and $H_{2021}$ is the set of highly-cited papers in 2021 computed as follows:
$H_{2021}$ $=\{i\in\{1,\dots,n\}\mid PUB(i)\in\{2019,2020\},$
@@
$\textit{NumArt}(w)=\#\{i\in\{1,\dots,n\}\mid w\in A_{i},\ 2020\leq\textit{PUB}(i)\leq 2022\}.$ (16)
@@
The quality of the journal is measured through the ISI impact factor, which is computed as the citation counts to papers the journal published in 2019-2020 from papers published in the year 2021 in other journals with an impact factor divided by the number of papers published by the journal in the years 2019-2020. As a reference for a high-impact prediction, we use the median impact factor among the top 10% journals in the “Health sciences” category (OECD category FORD 3030, data for the year 2021). To represent the research reputation of the authors, we use the maximum of the first and last authors’ Hirsch-index (H-index) values. The H-index is defined as the number of papers published by the author that have at least as many citations as this number. As a reference for a high-impact prediction, we use the median H-index of full professors, which was determined by analyzing the medical education journal editorial boards [26]. Previous research has found that the H-index was a good indicator of citation in the first few years after publication [25].
We also include additional impact metrics from social media as retrieved from Altmetric.com in September 2022: the number of tweets mentioning the research and the aggregate attention received on social media (such as the number of Facebook likes) weighted by the relative reach of each type of source (Facebook, YouTube, Twitter).
@@
Figure 4: This figure shows the top papers by predicted impact (Score) published in 2021. The reference impact factor (blue line) represents the median among the top 10% journals in the category “Health sciences” category. The reference H-index (red line) corresponds to the median H-index of full professors. The attention score and tweets were retrieved from Altmetric.com.
@@ Discussion
In Table 1(d), we also considered similar and important features related to the drug remdesivir, which has been used as a diagnostic word to assess word similarities for COVID-19 literature. In this case, all of the top important features were small-molecule COVID-19 drugs, with one exception being the antibody drug sarilumab approved for rheumatoid arthritis and repurposed for COVID-19. We note again that the suffix -ir, indicating an antiviral compound, is found for 8 of the 14 most important words similar to remdesivir, suggesting that subtoken approaches may provide additional benefits. Included in these top hits are some synonyms, e.g. molnupiravir and eidd-2801 (a preclinical name for this drug), lopinavir-ritonavir as well as lopinavir/ritonavir and lpv/r. There are 11/14 instances of first mentions in the literature prior to 2020, all representing drugs that were repurposed for COVID-19. Thus, we note that with our methodology, it is also possible to automatically identify instances of drug repurposing from the literature.
@@ Conclusions
The work of M. Černý was supported by Czech Science Foundation under project 25-18028S. The work of L. Beranová Dvořáčková was supported by grant IGA 16/2022 “PRECOG: Predicting REsearch COncepts of siGnificance”. M.P. Joachimiak was supported by the Director, Office of Science, Office of Basic Energy Sciences, of the US Department of Energy, Contract No. DE-AC0205CH11231. T. Kliegr was supported by the long-term support for research activities of FIS VSE and the EU under agreement No 857446 (HeartBit 4.0). The authors would like to thank Karel Douda for acquiring the citation counts.